OpenStudy (anonymous):

find the radius of the circle with these three points (-3-5),(-9,-5),(-6,-2) and center (-6,-5)

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

find distance from center to any one of the 3 points on circle... lgba.. i think this is 2D only

OpenStudy (lgbasallote):

lol yeah =)) sorry

OpenStudy (hba):

i guess the 3rd point gives a diffrent answer (-6,-2)

ganeshie8 (ganeshie8):

you can use the distance forumula : \(\sqrt{(x2-x1)^2+(y2-y1)^2}\)

OpenStudy (hba):

yeah thats right with any of the point

ganeshie8 (ganeshie8):

yeah @Colocha can you find the radius using distance formula ?

OpenStudy (anonymous):

um i gt 2.449

OpenStudy (anonymous):

its probably wrng

OpenStudy (anonymous):

o i messed up

ganeshie8 (ganeshie8):

hmm which two points you took... .one point must have to be center, other can be any of the three

OpenStudy (anonymous):

o ok ill do it again

ganeshie8 (ganeshie8):

np good luck :) you should get 3 maybe

OpenStudy (hba):

actually u have to take the radius and any one point along with it

OpenStudy (anonymous):

i gt three i choose (-9,-5) and the center

OpenStudy (hba):

i mean the center

OpenStudy (hba):

yeah then ur right good job

OpenStudy (anonymous):

ook now this one find the radius of the circle defined by the equation x^+y^2-10x+4y+4=0

OpenStudy (hba):

please post it in other link

ganeshie8 (ganeshie8):

okay distance between (-6, -5) and (-9, -5) = \(\sqrt{(-6-(-9))^2+(-5-(-5)^2)}\) \(\sqrt{(-6 + 9)^2+(-5 + 5)^2}\) \(\sqrt{(3)^2+(0)^2}\) = ?

ganeshie8 (ganeshie8):

to find the radius of x^+y^2-10x+4y+4=0 ---------(1) rearrange the equation in standard form : (x-h)^2 + (y-k)^2 = r^2 ---------------------(2) "r" is the radius

OpenStudy (anonymous):

idk how to do that

ganeshie8 (ganeshie8):

Ok. its somewhat tricky to arrange the equation in standard form. pls post it in a new thread... il see if i can help :))

OpenStudy (anonymous):

its on a new thread